# A Non-Traditional Approach to Assisting Data Address Translation

### v1

@@ IV-E PCAT Indexing
Figure 12 presents the CDTLB miss rates for cache block address and load PC indexing of a PCAT. The miss rates are very similar in most cases, with either performing slightly better in some cases. We use the block address indexing for all the results we present subsequently because it results in only a single PCAT access for a block of instructions.
IV-F Reduction in Primary CDTLB Misses
Figure 13: Effective Load Miss Rate (64-Entry CDTLB)
Figure 14: Effective Load Miss Rate (128-Entry CDTLB)
Figure 15: Comparing 128-entry CDTLB with PCAX
@@ IV-F Reduction in Primary CDTLB Misses
We now consider the reduction in the primary CDTLB misses due to the translation requests fulfilled by a PCAT1. Figures 13 and 14 show the effective CDTLB miss rate for 64- and 128-entry, 4-way set associative CDTLBs, respectively. The effective miss rate is the miss rate for load references that can’t be serviced by the PCAT1 and CDTLB, and end up accessing the STLB. In each case, we plot the base CDTLB miss rate (Base) and the effective miss rate with 32-, 64-, 128-entry, 2-way set associative PCAT1s. The PCAT2 size is 2048 entries and is 8-way set associative in all cases.
The figures show a significant reduction in the effective CDTLB miss rate in almost all cases, even with a 32-entry PCAT1. A 64-entry PCAT1 can achieve more reduction in several cases; reduction beyond 64-entries is marginal. The magnitude of the reduction is quite impressive. From figure 13 observe that while the base miss rate is 4%-12% for most of the applications, it is brought down to 2%-5% with a 64-entry PCAT1. For many benchmarks, there is a 3X reduction in the effective miss rate. The significant improvements continue even for a 128-entry CDTLB in figure 14. There is a greater than 2X reduction in the miss rate with a PCAT, with base miss rates in the 4%-6% range brought down to the 2%-2.5% range for many applications.
@@
Figure 15 combines select information presented in Figures 13 and 14 to allow for a direct comparison between a 64-entry CDTLB, a 128-entry CDTLB, and a 64-entry CDTLB and 64-entry PCAT1 (PCAX). The latter two take up essentially the same amount of storage. As we can see, PCAX outperforms the 128-entry CDTLB in reducing the miss rate in almost all cases. Later in sections IV-J and IV-K we will also see that PCAX gets better performance and consumes less energy, respectively, than the larger 128-entry CDTLB.
@@
Next, we consider the impact of PCAT2 size in Figure 16. Here we have a 64-entry, 4-way set associative CDTLB and a 64-entry, 2-way set associative PCAT1. We have plots for the effective CDTLB load miss rate for PCAT2 configurations of 512, 1K, 2K, and 4K, with 8-way set associativity in all cases, in addition to the base CDTLB (Base) and also simulate a configuration with no PCAT2. PCAT2 sizes of 512 and 1K entries do quite well for many applications, but leave some opportunity for improvement in others. A 2K entry PCAT2 is nearly as effective as the ones with 4K entries. Referring back to Figure 7, barring conflicts due to limited associativity, given that most applications had fewer than 2000 problem loads, we expected a 2K entry PCAT2 to be quite effective. For many applications that had fewer than 500 problem loads, a 512-entry PCAT2 would be adequate. The results of Figure 16 bear this out. Not using a PCAT2 does not achieve the benefits of PCAX in several cases. This is because the PCAT entry is lost when the load is evicted from the L1i and thus is not available for reuse when the load is brought back into the L1i for processing.
IV-G Large Page Sizes
@@
Figure 18: Large Page Improvements with PCAX
Figure 19: Large Page Comparison with 4K+PCAX
We now consider the effectiveness of PCAX with (uniform) page sizes larger than 4KB. We consider only a uniform page size as our traces do not have multiple page sizes. Figure 17 presents the percentage of loads that miss in a 64-entry, 4-way CDTLB that accesses the same PTE as the previous dynamic instance. We consider page sizes of 4KB, 16KB, 64KB, 256KB, and 1MB, and present data for benchmarks where the CDTLB miss rate is greater than 1%. Clearly, the phenomenon that underlies PCAX—that a large number of misses access the same PTE as the previous dynamic instance—is prevalent even for larger pages. Figure 18 presents the miss rate improvement for larger page sizes using PCAX. The CDTLB miss rate improvements continue even for larger page sizes.
Figure 19 presents CDTLB miss rates for large pages and also for 4KB pages with PCAX. Clearly, using larger page sizes brings down the miss rate. More interestingly using 4K pages with PCAX provides lower miss rates compared to 16K and 64K page sizes. It performs as well for many benchmarks or comes close to the performance using a 256K page size. This suggests that with PCAX larger page sizes may not be needed if a reduction in CDTLB misses is the objective.
Figure 20: Effective STLB Miss Rate
@@
IV-H Reduction in STLB Misses
Because the PCATs are not impacted by the normal operation and replacements in the CDTLB and STLB, they are able to hold on to address translations that would miss in them for a longer period. Further, by not directly placing a PTE in the STLB on an STLB miss, and placing it in the PCAT1 and CDTLB instead, allows other (potentially useful) PTEs to continue to reside in the STLB, rather than be evicted. This helps reduce the number of STLB misses. Figure 20 shows the effective STLB load (global) miss rate for the base case, and for PCAT2 sizes of 512, 1K, 2K, 4K, and 8K entries, with 8-way set associativity in each case. The PCAT1 is a 64-entry, 2-way set associative. To improve readability, we only plot the applications where the Base miss rate is greater than 1%. With PCATs, we can achieve a noticeable reduction in the effective STLB (global) miss rate. Though the absolute magnitude is small, the percentage reduction is meaningful, especially as a miss in the STLB results in expensive page table accesses (Page Walks), which can involve several memory references.
@@ IV-I Data Address Translation Service Sites
We now consider the different sites from where a data address translation is serviced in Figure 21. The applications in the figure represent a subset of the applications merely to improve readability, as presenting data for all 100 applications would be cluttered. Applications not considered show similar results, and we will consider all the applications when we factor this data into the CPI impact in section IV-J and energy impact in section IV-K. The results are for a 64-entry, 4-way CDTLB, 1536-entry, 12-way STLB, 64-entry, 2-way PCAT1, a 2K entry, 8-way PCAT2, and a 2-entry RPA buffer.
With conventional address translation, a valid PTE can be obtained from (i) the CDTLB (labeled CDTLB), (ii) the STLB in case of a CDTLB miss and STLB hit (labeled STLB), or (iii) the Page Table in case of an STLB miss (labeled Page Walk). The left bar for a benchmark in the figure shows the percentage of load references that get a PTE from these different sites. With a PCAT, a valid PTE could additionally be obtained from the PCAT in three different situations: (a) the load instruction that was used to access the PCAT1 and the translation access would have missed in the CDTLB (labeled PCAT-M), (b) the same load instruction and the translation access would have hit in the CDTLB (labeled PCAT-HS), or (c) a different load instruction whose translation access would have hit in the CDTLB (labeled PCAT-HD). For (b) and (c) the PTE is obtained from the RPA buffer. The right bar for a benchmark shows the percent of load instructions that get their address translation in each of the six different ways. The Y-axis uses a log scale so that the difference in the important portions, PCAT-M and Page Walk, can be better appreciated.
Observe that the number of references served by the CDTLB and the STLB are cut down significantly in most cases because of the references served by the PCAT (PCAT-M, PCAT-HS, and PCAT-HD). There is even a non-trivial reduction in expensive Page Walks in cases where STLB misses are non-negligible. These reductions in CDTLB and STLB accesses, and Page Walks, lead to a lower data translation CPI and energy as we quantify next.
@@ IV-J CPI for Data Address Translation
In this section, we consider the improvement in the CPI for load data address translation with PCAX, borrowing from the CPI analysis in [26], and the potential performance benefit. The parameters of the CPI model are shown in Table II. Our computation of cycles spent in load address translation in conventional components consists of: (i) Cycles ($Cycles_{L1-H}$) spent in servicing CDTLB Hits ($H_{L1-H}$), where hit latency is 1 cycle (ii) Cycles ($Cycles_{L1-M}$) spent servicing CDTLB Misses ($M_{L1-M}$), which lookup the STLB that takes 8 cycles (iii) Cycles ($Cycles_{L2-M}$) spent servicing STLB misses ($M_{L2-M}$) which takes an average of 30 cycles. Since PCAX provides a PTE before the address calculation, we consider it to take 0 cycles for address translation.
Figure 22 shows the data address translation CPI for load instructions for a 64-entry, 4-way CDTLB (Base-64) and a 128-entry, 4-way CDTLB (Base-128) and a 64-entry, 4-way CDTLB with PCAX with default PCAT1/PCAT2 sizes (PCAX) for the benchmarks. PCAX achieves a lower effective CPI, with the 64-entry CDTLB and 64-entry PCAT1 configuration having a lower CPI than even a 128-entry CDTLB in most cases.
The lower data translation CPI also results in improved overall performance. Figure 23 shows the percentage improvement in performance (over Base-64), obtained from a timing simulation that incorporates a detailed TLB hierarchy and page walk modeling (as was used in [50, 49])666 The ordering of the benchmarks on the X-axis in Figure 23 is different from that in Figure 22.. The performance improvement is on par with many other microarchitectural techniques. For most benchmarks (except 5 of 84), the 64-entry CDTLB with PCAX configuration has better performance improvement than a 128-entry CDTLB (even though we optimistically model the same access latency for both CDTLB sizes).
@@
Figure 22: CPI per load with PCAX
@@
Figure 24: Energy improvement with PCAX
@@ IV-K Energy Impact
Next, we consider the energy impact of PCAX over a base 64-entry, 4-way CDTLB. We measure energy consumption using CACTI 7.0 [29] with 22nm technology. The dynamic energy per Read and Write Operation for the conventional address translation structures and the PCATs are shown in Table III. It is worth noting that the energies of the PCAT1 and PCAT2 are smaller than the CDTLB and STLB because they can be run a lot slower. In the case of PCAT1, it can take multiple cycles (5 in our case) to respond and still provide a translation before the address calculation. Likewise, PCAT2 (10 cycles) is accessed alongside the (slower) L2 cache. The longer latencies also allow for sequential tag and data access for energy savings, which we use for PCAT1 and PCAT2.
Figure 24 shows the impact of PCAX on the dynamic energy for data address translation 777 The ordering of the benchmarks on the X-axis in Figure 24 is different from the ordering in Figures 22 and 23.. The numbers in the figure account for both loads and stores. If we considered the energy for only loads, the energy benefit would be higher as stores have little energy benefit. The figure shows a small energy cost in a few benchmarks. These are also the cases that have a smaller number of CDTLB and STLB misses, to begin with, and PCAX provides a smaller reduction in the absolute number of misses. However, for most benchmarks, there is a meaningful improvement in dynamic energy, in addition to the CPI and performance benefits we saw earlier. The energy reduction is higher in cases when page walks are reduced due to fewer STLB misses, resulting in fewer L1D cache accesses due to fewer page walks. A 128-entry, 4-way CDTLB (Base-128), on the other hand, takes more energy than a 64-entry CDTLB (improvement is negative) in most cases.
@@ V Related Work
Given the importance of data address translation, there has been a plethora of work in improving the efficiency of the process. The overarching theme of the vast body of work is to decrease the number of DTLB “misses”, i.e., the number of times an address translation requires accessing the secondary TLBs or the page tables. We provide a brief overview of some of this prior work below; excellent summaries of such work can also be found in some recent papers [21, 50, 22, 26].
There is a large body of work on novel DTLB designs. Some of these designs [37, 35, 45] seek to exploit contiguity in both the virtual and physical pages and minimize the number of TLB entries, thereby reducing the DTLB misses. Processors and the OS provide support [2] for large pages to increase the TLB reach and reduce the number of DTLB misses. Different TLB designs have been explored to better support huge pages [15, 43, 33, 26]. Huge pages rely on memory contiguity, which might be absent due to fragmentation and a plethora of techniques have been proposed to work around this concern [9, 34, 25, 19, 32, 44, 46]. Huge pages exploit discrete levels of contiguity, [22] is a scheme that allows exploiting intermediate levels of contiguity. Techniques to reduce TLB misses have also employed prefetching (primarily into the secondary TLB), exploiting memory access stream patterns [11, 24, 42]. A paper [50] proposes combining prefetchers and picking one of multiple depending on their usefulness and exploits locality in the page table to cache nearby PTEs and minimize page walks (and effectively prefetching PTEs). Another recent paper to increase TLB reach involves reorganizing the page table and using hash functions to significantly reduce miss rates for data-intensive applications [21] .
Techniques have also been proposed to reduce the TLB miss cost (mainly page walk cost) by caching parts of the page walk in the memory hierarchy. Such techniques employ MMU caches to cache higher levels of the page table and reduce the latency of a page walk [8, 10, 40, 1], and are found in many modern processors. PTEs are also cached to reduce page walk latency.
@@
The VAX processors used the concept of pre-translation, which observed that multiple dynamic instances of instruction fetching accessed the same PTE and optimized it by maintaining the PTE in a special register [41]. Regarding optimizations to the processing of load instructions based upon their identity rather than the data address, researchers have also made an empirical observation that a small number of static load instructions account for a disproportionate number of data cache misses [4, 51, 48] and proposed appropriate optimizations. Instruction PC-based prediction has also been proposed to optimize coherence events due to store and load instructions in a multiprocessor [27].

### v2

@@ IV-F PCAT Indexing
Figure 13 presents the CDTLB MPKIs for cache block address and load PC indexing of a PCAT. The MPKIs are very similar in most cases, with either performing slightly better in some cases. We use the block address indexing for all the results we present subsequently because it results in only a single PCAT access for a block of instructions.
IV-G Reduction in Primary CDTLB Misses
We now consider the reduction in the primary CDTLB misses due to the translation requests fulfilled by a PCAT1. Figures 14 and 15 show the effective CDTLB MPKI for 64- and 128-entry, 4-way set associative CDTLBs, respectively. The effective MPKI is the MPKI for load references that can’t be serviced by the PCAT1 and CDTLB, and end up accessing the STLB. In each case, we plot the base CDTLB MPKI (Base) and the effective MPKI with 32-, 64-, 128-entry, 2-way set associative PCAT1s. The PCAT2 size is 2048 entries and is 8-way set associative in all cases. Detailed data for each benchmark is shown in the figures for the above PCAT1 configurations but below we also report geometric mean (GM) averages for 8-way and fully-associative configurations for which detailed data is not presented in the figures.
Figure 14: Effective Load MPKI (64-Entry CDTLB)
Figure 15: Effective Load MPKI (128-Entry CDTLB)
Figure 16: Comparing Base-Iso with PCAX
@@ IV-G Reduction in Primary CDTLB Misses
The figures demonstrate a substantial reduction in the effective CDTLB MPKI across nearly all workloads, even with a modest 32-entry PCAT1. Increasing the PCAT1 to 64 entries provides additional reductions for several applications, although benefits beyond 64 entries are marginal. The magnitude of improvement is noteworthy. From Figure 14, observe that while the baseline MPKI ranges from 5-13 (geometric mean (GM): 8.19) for most applications, a 64-entry PCAT1 lowers this to 2-5 (GM: 3.43). For 10 benchmark traces, the effective MPKI is reduced by approximately 3X, and for 68 traces it is reduced by more than 2X. Similar improvements occur with an 8-way associative CDTLB: the baseline CDTLB MPKI GM of 6.75 is reduced to 2.64 with PCAX. With a fully associative CDTLB, the CDTLB MPKI GM decreases from 4.85 (baseline) to 2.23 with PCAX. Significant gains persist even with a 128-entry CDTLB, as shown in Figure 15. For 67 benchmark traces, the MPKI reduction exceeds 2X, with the baseline MPKI in the 4-7 range (GM: 4.45) for most applications reduced to 2-3 (GM: 1.77). The improvements with an 8-way associative CDTLB follow the same trend: the baseline CDTLB MPKI GM of 4.45 drops to 1.42 with PCAX. Under fully associative designs, the CDTLB MPKI GM decreases from 2.01 (baseline) to 0.99 with PCAX.
@@
Figure 16 combines select information presented in Figures 14 and 15 to allow for a direct comparison between a 64-entry CDTLB, a 128-entry CDTLB and a 4K-entry STLB (Base-Iso), and a 64-entry CDTLB and 64-entry PCAT1 (PCAX). The latter two take up essentially the same amount of storage, and hence we refer to this configuration as Base-Iso in the coming sections. As we can see, PCAX outperforms the Base-Iso configuration in reducing the MPKI in almost many cases. Later in sections IV-J and IV-K we will also see that PCAX gets better performance and consumes less energy, respectively, than Base-Iso. Across workloads, the GM CDTLB MPKI is 8.2 for Base-64, 4.45 for Base-Iso, and 3.4 for PCAX.
@@
Next, we consider the impact of PCAT2 size in Figure 17. Here we have a 64-entry, 4-way set associative CDTLB and a 64-entry, 2-way set associative PCAT1. We have plots for the effective CDTLB load MPKI for PCAT2 configurations of 512, 1K, 2K, and 4K, with 8-way set associativity in all cases, in addition to the base CDTLB (Base-64) and also simulate a configuration with no PCAT2. PCAT2 sizes of 512 and 1K entries do quite well for many applications, but leave some opportunity for improvement in others. A 2K entry PCAT2 is nearly as effective as the ones with 4K entries. Referring back to Figure 11, barring conflicts due to limited associativity, given that most applications had fewer than 2000 problem loads, we expected a 2K entry PCAT2 to be quite effective. For many applications that had fewer than 500 problem loads, a 512-entry PCAT2 would be adequate. The results of Figure 17 bear this out. Not using a PCAT2 does not achieve the benefits of PCAX in several cases. This is because the PCAT entry is lost when the load is evicted from the L1i and thus is not available for reuse when the load is brought back into the L1i for processing. The GM MPKI is 8.2 for Base-64, 6.0 without PCAT2, 4.2 with a 512-entry structure, 3.6 with 1K entries, and 3.4 with both 2K and 4K entries.
IV-H Large Page Sizes
We now consider the effectiveness of PCAX with (uniform) page sizes larger than 4KB. We consider only a uniform page size as our traces do not have multiple page sizes. Figure 18 presents the percentage of loads that miss in a 64-entry, 4-way CDTLB that accesses the same PTE as the previous dynamic instance. We consider page sizes of 4KB, 16KB, 64KB, 256KB, and 1MB, and present data for benchmarks where the CDTLB miss rate is greater than 1%. Clearly, the underlying phenomenon—that a large number of misses access the same PTE as the previous dynamic instance—is prevalent even for larger pages. Figure 19 presents the MPKI improvement for larger page sizes using PCAX. The CDTLB MPKI improvements continue even for larger page sizes. Note that though the MPKI reductions might be higher for larger page sizes in some cases, the base miss rates are also smaller for the large page sizes. The key takeaway is that PCAX continues to provide benefit at bigger page sizes.
@@ IV-H Large Page Sizes
Figure 19: % MPKI Reductions with PCAX
Figure 20: Large Page MPKI Comparison with 4K+PCAX
Figure 20 presents CDTLB MPKIs for large pages and also for 4KB pages with PCAX. Clearly, using larger page sizes brings down the MPKI. More interestingly using 4K pages with PCAX provides lower MPKIs compared to 16K and 64K page sizes. It performs as well for many benchmarks or comes close to the performance using a 256K page size. This suggests that with PCAX larger page sizes may not be needed if a reduction in CDTLB misses is the objective.
@@ IV-I Data Address Translation Service Sites
We now consider the different sites from where a data address translation is serviced in Figure 21. Data in the figure is for a subset of the benchmarks to improve readability, as presenting data for all 84 applications would be cluttered. Applications not considered show similar results, and we will consider all the applications when we factor this data into the CPI impact in section IV-J and energy impact in section IV-K. For each benchmark we have three stacked bars: (i) the left bar is for a conventional 64-entry, 4-way CDTLB, 1536-entry, 12-way STLB, (ii) the middle bar is the the Base-Iso configutation (128-entry, 4-way CDTLB, 4K-entry, 16-way STLB), and (iii) the right bar is for PCAX (64-entry, 4-way CDTLB, 1536-entry, 12-way STLB, 64-entry, 2-way PCAT1, a 2K entry, 8-way PCAT2, and a 2-entry RPA buffer).
With conventional address translation, a valid PTE can be obtained from (i) the CDTLB (labeled CDTLB), (ii) the STLB for a CDTLB miss and STLB hit (labeled STLB), or (iii) the Page Table for an STLB miss (labeled Page Walk). With a PCAT, a valid PTE could additionally be obtained from the PCAT in three different situations: (a) the load instruction that was used to access the PCAT1 and the translation access would have missed in the CDTLB (labeled PCAT-M), (b) the same load instruction and the translation access would have hit in the CDTLB (labeled PCAT-HS), or (c) a different load instruction whose translation access would have hit in the CDTLB (labeled PCAT-HD). For (b) and (c) the PTE is obtained from the RPA buffer where it was placed after a PCAT1 access. Each bar for a benchmark shows the percent of load instructions that get their address translation in each of the six different ways. The Y-axis uses a log scale so that the difference in the important portions, PCAT-M and Page Walk, can be better appreciated.
Observe that the number of references served by the CDTLB significantly in most cases because of the references served by the PCAT (PCAT-M, PCAT-HS, and PCAT-HD). There is even a non-trivial reduction in expensive Page Walks in some cases where STLB misses are non-negligible. Note that references PCAT-HS and PCAT-HD do not yield in any CPI or IPC benefit, they help mainly with energy reduction. These reductions in CDTLB and STLB accesses, and Page Walks, lead to a lower data translation CPI and energy as we quantify next.
@@ IV-J CPI for Data Address Translation
In this section, we consider the improvement in the CPI for load data address translation with PCAX, borrowing from the CPI analysis in [22], and the potential performance benefit. The parameters of the CPI model are shown in Table II. Our computation of cycles spent in load address translation in conventional components consists of: (i) Cycles ($Cycles_{L1-H}$) spent in servicing CDTLB Hits ($H_{L1-H}$), where hit latency is 1 cycle (ii) Cycles ($Cycles_{L1-M}$) spent servicing CDTLB Misses ($M_{L1-M}$), which lookup the STLB that takes 8 cycles (iii) Cycles ($Cycles_{L2-M}$) spent servicing STLB misses ($M_{L2-M}$) which takes an average of 30 cycles.
Figure 22 shows the data address translation CPI for load instructions for a 64-entry, 4-way CDTLB (Base-64), Base-Iso, and a 64-entry, 4-way CDTLB with PCAX with default PCAT1/PCAT2 sizes (PCAX) for the benchmarks. PCAX achieves a lower CPI, with the 64-entry CDTLB and 64-entry PCAT1 configuration having a lower CPI than even Base-Iso in most cases. The GM CPI is 1.52 for Base-64, compared to 1.26 for Base-Iso and 1.19 for PCAX.
The lower data translation CPI also results in improved overall performance. Figure 23 shows the percentage improvement in performance (over Base-64), obtained from a timing simulation that incorporates a detailed TLB hierarchy and page walk modeling (as was used in [40, 39]). Note: The benchmark ordering on the X-axis in Figure 23 differs from that in Figure 22. The performance improvement is on par with many other microarchitectural techniques. GM performance improvement with PCAX is 1.7% over the base configuration and 1% for the Base-Iso configuration over the base configuration. For most benchmarks (except 6 of 84), the 64-entry CDTLB with PCAX configuration has better performance improvement than Base-Iso (even though we optimistically model the same access latency for both CDTLB/STLB sizes).
@@
Figure 22: Data address translation CPI per load with PCAX
@@
Figure 24: Energy improvement (%) with PCAX
@@ IV-K Energy Impact
Next, we consider the energy impact of PCAX over a base 64-entry, 4-way CDTLB. We measure energy consumption using CACTI 7.0 [24] with 22nm technology. The dynamic energy per Read and Write Operation for the conventional address translation structures and the PCATs are shown in Table III. It is worth noting that the energies of the PCAT1 and PCAT2 are smaller than the CDTLB and STLB because they can be run a lot slower. In the case of PCAT1, it can take multiple cycles (5 in our case) to respond and still provide a translation before the address calculation. Likewise, PCAT2 (10 cycles) is accessed alongside the (slower) L2 cache. The longer latencies also allow for sequential tag and data access for energy savings, which we use for PCAT1 and PCAT2. The energy calculation accounts for the use of access PCAT bits, which determine the number of accesses to PCATs.
Figure 24 shows the impact of PCAX on the dynamic data address translation energy; the benchmark ordering on the X-axis in Figure 24 differs from that in Figures 22 and 23. The numbers in the figure account for both loads and stores. For only loads, the energy benefit is higher as stores have little energy benefit. There is a meaningful improvement in dynamic energy (in addition to the CPI and performance benefits seen earlier) in most cases. The energy reduction is higher in cases when page walks are reduced due to fewer STLB misses, resulting in fewer L1D cache accesses due to fewer page walks, with a GM energy reduction of 7%. The Base-Iso configuration on the other hand, takes more energy than a 64-entry CDTLB (improvement is negative) in most cases, with a GM increase in energy consumption of 8.5%.
@@ IV-L Effect of Invalidating PCATs
IV-M Analysis on other workloads
As mentioned above, the benchmarks we used in this paper have have been widely used in recent papers using the Champsim simulation infrastructure, and studying a variety of issues in instruction supply, memory hierarchies, and microarchitecture. They represent a class of server workloads with large code footprints, and in many of them data address translation is also an important problem. To assess the prevalence of the observations in section , we briefly examined the translation behavior across several additional workload classes, including Google datacenter traces [31], SPEC applications, and graph workloads.
For the Google datacenter traces for the x86 architecture, we observe similar qualitative behavior. In most traces, more than 50% of the CDTLB misses occur when a load misses on the same PTE referenced by its previous dynamic instance. Although the overall CDTLB miss rates in these traces are generally lower than those observed in our primary workloads, approximately 50–55% of the misses still exhibit this per-PC PTE recurrence.
For SPEC applications with CDTLB miss rates greater than 1%, we observe strong per-PC PTE recurrence for cactuBSSN and gamess (over 90%), and moderate recurrence for gcc (about 72%). The effect is weaker for omnetpp and gromacs (around 25%), and largely absent for mcf, GemsFDTD, and astar, which do not exhibit this regularity. Many other SPEC and other integer and floating point applications had a very small CDTLB miss rate.
Finally, we briefly examined graph workloads, which are often highly translation-bound due to their extremely large data footprints. Although these applications typically have a small number of static load PCs responsible for most CDTLB misses, those load PCs do not repeatedly access the same virtual page across dynamic instances (similar to mcf). Instead, their pointer-chasing behavior causes a given static load PC to reference many different pages over time, breaking the PC→PTE stability that PCAX relies upon. As a result, despite their high DTLB MPKI, graph workloads derive limited benefit from PCAX.
Overall, these observations suggest that PCAX is most effective for workloads that simultaneously exhibit: (i) non-trivial CDTLB miss rates, (ii) large instruction working sets that create temporal separation between dynamic instances of load PCs, and (iii) structured data-access patterns that preserve per-PC virtual page locality.
@@ V Related Work
Given the importance of data address translation, a large body of work aims to reduce DTLB misses, i.e., DTLB accesses that fall back to secondary TLBs or page tables. Excellent overviews appear in recent papers [17, 40, 18, 22] and we provide a brief overview of some of the prior work below.
A significant line of work proposes novel DTLB designs. Several designs exploit virtual and physical page contiguity to reduce the number of TLB entries and misses [30, 28]. Hardware and OS support for large pages [37] similarly improves TLB reach, and many designs tailor the TLB to better support huge pages [13, 26, 22]. Since contiguity may be limited by fragmentation, techniques have also been proposed to mitigate this challenge [5, 27, 21]. Beyond discrete large-page sizes, [18] exploits intermediate levels of contiguity. Other recent work reorganizes page tables and employs hashing to significantly reduce miss rates [17]. Several proposals cache page-table components (e.g., via MMU and PTE caches) to reduce TLB misses and page-walk latency [4, 7, 33, 19].
Another thread of work reduces misses through prefetching—primarily into secondary TLBs—by exploiting memory access stream patterns [6, 20, 35]. Recent work combines multiple prefetchers and selects between them based on usefulness [40], also leveraging locality in the page table to prefetch nearby PTEs.
PC‑based TLB prefetchers ([20, 39]) appear superficially related to PCAX because both use the program counter (PC) as a correlating signal. However, their goals, mechanisms, and semantics differ fundamentally. PC‑based prefetchers use the PC to predict future virtual page numbers and speculatively prefetch those pages into the STLB. The PC serves only as a trigger to initiate speculative requests. PCAX does not predict future VPNs: instead, it stores and supplies the actual PTE most recently missed by that static load PC. Rather than prefetching into the TLB hierarchy, PCAX provides an alternative translation path, bypassing the CDTLB entirely for select loads by returning a PTE at fetch time. As a result, PCAX is not a prefetching mechanism; it is a PC‑driven address‑translation path that directly supplies the PTE when the PC reappears.
In contrast to memory‑address‑stream speculation, PCAX exploits the stability of translation behavior for specific load PCs, binding translations directly to instruction identity rather than virtual‑address patterns. This enables PCAX to access the PTE during instruction fetch, well before the data address is computed or the CDTLB is accessed. Because the PTE is not needed until the address‑generation stage, PCAX can perform its lookup slowly and energy‑efficiently without impacting performance. This early, PC‑driven mechanism provides PTEs for select loads ahead of time, improving both performance and data‑translation energy, and operates orthogonally to both STLB replacement and prior prefetching techniques.
The VAX processors used the concept of pre-translation, which observed that multiple dynamic instances of instruction fetching accessed the same PTE and optimized it by maintaining the PTE in a special register [34]. In a similar vein, our design incorporates an RPA buffer that opportunistically reuses recently supplied translations from PCAX.
@@
Regarding optimizations to the processing of load instructions based upon their identity rather than data address, researchers have also made empirical observations that a small number of static load instructions account for a disproportionate number of data cache misses [1, 38].